# ImAgent: A Unified Multimodal Agent Framework for Test-Time Scalable Image Generation

### v2

@@ 1 Introduction
Text-to-Image (T2I) models [rombach2022high, ramesh2022hierarchical, podell2023sdxl, sun2024autoregressive, tian2024visual, qin2025lumina] have made remarkable progress in generating visually realistic and semantically coherent images from natural language descriptions. Despite these advancements, the quality of generated results often exhibits randomness and inconsistency with the given prompts [xu2024prompt, lee2024direct]. This limitation primarily stems from the strong dependency of existing T2I models on the clarity and specificity of textual inputs. When the prompt is vague or underspecified, the models tend to overlook key semantic elements, leading to images that deviate from the intended meaning and fail to fully capture the user’s intent.
To mitigate this problem, prior research has proposed various strategies, such as prompt rewriting [brade2023promptify, mo2024dynamic, liu2025one], best-of-N sampling [li2025reflect, singhal2025general], classifier-free guidance [ho2022classifier, sauer2023stylegan, shen2024rethinking], and self-revision or iterative refinement [singh2023divide, li2024g, jeon2025iterative]. These approaches aim to either clarify the textual prompt or reduce the model’s inherent randomness rather than finetuning the models, reflecting the underlying principle of test-time scaling—improving generation quality through additional inference-time computation.
@@
In this paper, we aim to construct a universal agent for image generation that can adaptively select the optimal action for a given case, allocate computational resources accordingly, and execute the selected action within the agent itself without relying on any external models. This design enables more efficient test-time scaling. Thanks to recent advances in unified multimodal models [deng2025emerging, wu2025janus, wu2025janus, xie2024show, qin2025lumina], which integrate text generation, image generation, and visual understanding within a single framework, this assumption becomes attainable. Building upon this foundation, we introduce ImAgent, a training-free unified multimodal agent designed to perform efficient test-time scaling for image generation.
@@ 2.1 Text-to-Image Generation
Text-to-Image (T2I) models have demonstrated remarkable ability in synthesizing visually realistic and semantically coherent images from natural language descriptions. The evolution of T2I models has followed several major paradigms, including diffusion-based models [podell2023sdxl, rombach2022high, ramesh2022hierarchical, esser2024scaling, gao2024lumina, xie2025sana], autoregressive models [sun2024autoregressive, zhang2024var, tian2024visual, wang2024emu3, zheng2025parallel, xiong2025llava], and more recently, unified multimodal models [xiao2025omnigen, xie2024show, zhou2024transfusion, tong2025metamorph, deng2025emerging, zheng2025learning, luo2024stable]. These approaches have significantly advanced image fidelity, diversity, and text-image alignment.
@@ 2.2 Unified Multimodal Models
Unified multimodal models refer to models that jointly possess multimodal understanding and generation capabilities within a single framework. Unlike traditional pipelines that rely on separate models for text or image understanding, image generation, and image editing, these models integrate vision and language processing into a shared architecture, enabling seamless information exchange across modalities [xiao2025omnigen, zhou2024transfusion, tong2025metamorph, wu2025janus, wu2025janus]. This convergence of understanding and generation not only improves efficiency and coherence but also paves the way for building more general-purpose, human-like agents capable of reasoning, creating, and self-correcting.
@@ 2.3 Generation Optimization Strategies
Despite the remarkable progress of T2I models, the quality of generated images still depends heavily on the clarity of prompts and the randomness inherent in the generation process. To alleviate these issues, various optimization strategies have been proposed. Best-of-N sampling [tian2025unigen, verdun2025soft, li2025reflect] generates multiple candidates and selects the best one, reducing the influence of stochastic variations in model sampling. Prompt rewriting [tian2025unigen, wang2024promptcharm, yasunaga2022retrieval] reformulates or enriches the input text to provide more explicit guidance for the generator, improving semantic alignment and visual coherence. classifier-free guidance [ho2022classifier, sauer2023stylegan, shen2024rethinking] controls the conditioning strength to balance fidelity and diversity, and iterative self-refinement [yang2023idea2img, madaan2023self, singh2023divide, li2024g, jeon2025iterative] enables models to assess and revise their outputs through multiple rounds of generation. Although these techniques significantly enhance performance, they typically operate as independent modules with limited interaction between reasoning, generation, and evaluation—highlighting the need for unified multimodal frameworks that integrate these capabilities into a cohesive pipeline.
@@ 3 Method
We construct ImAgent upon unified multimodal models, which inherently possess both understanding and generation capabilities. Motivated by the observation that multimodal understanding is generally more reliable and easier to achieve than multimodal generation [zhang2025unified, zhang2025unified1], ImAgent leverages the model’s strong understanding ability to adaptively select the optimal action for a given case that enhances the overall generation quality. This design enables ImAgent to fully exploit the potential of a single unified multimodal model without any additional training or external modules, thereby achieving efficient test-time scaling. In this section, we first introduce the overall agent framework of ImAgent, followed by a detailed description of its action space.
@@ 3.1 Agent Framework Overview
This iterative process continues until the agent selects the STOP action, which indicates that the generated result is satisfactory, or until the maximum number of steps (set to 5 by default) is reached. Notably, both the policy controller and the action execution are implemented within a single unified multimodal model. The detailed algorithm for image generation is presented in Algorithm 1, while the algorithm for image editing follows a similar procedure and is provided in the Appendix.
@@
| Vanilla | 0.47 | 0.65 | 0.53 | 0.68 | 0.40 | 0.54 | |
| Bagel | ImAgent | 0.58 | 0.68 | 0.62 | 0.71 | 0.53 | 0.62 |
| Vanilla | 0.45 | 0.60 | 0.46 | 0.46 | 0.36 | 0.47 | |
| Janus-Pro-7B | ImAgent | 0.47 | 0.58 | 0.52 | 0.54 | 0.44 | 0.51 |
3.2 Action Space
In this section, we introduce the action space of ImAgent. The detailed definitions and usage of each action are provided in the Appendix.
Naive Generation/Editing.
Prompt Enhancement with CoT.
This action refines the input prompt by enriching vague or underspecified user queries with additional contextual and descriptive details. Since text-to-image (T2I) models are highly sensitive to prompt wording [brade2023promptify, mo2024dynamic, liu2025one], more elaborate and specific prompts typically lead to higher-quality visual outputs. However, most T2I backbones are trained on simple captions or surface-level image descriptions, which limits their ability to reason over complex or compositional instructions. To address this limitation, we leverage the language reasoning capability of the model’s understanding module through Chain-of-Thought (CoT) prompting. By explicitly performing intermediate reasoning and elaboration, the agent transfers structured linguistic insights from the understanding domain to the generation process, thus producing more semantically aligned and visually coherent outputs.
Prompt Revision Based on the Generated/Edited Image.
Image Detail Refinement.
This action aims to correct minor imperfections in the generated or edited image when the input instruction is already satisfactory and the remaining issues stem from the generation module itself. It refines local visual details—such as textures, lighting, or small artifacts—without modifying the current prompt. By enhancing fine-grained fidelity while maintaining semantic consistency, this action improves the overall perceptual quality and realism of the generated results.
Best-of-N Sampling.
This action mitigates the inherent stochasticity of text-to-image (T2I) generation. Unlike language models that typically produce stable, low-entropy outputs through supervised fine-tuning (SFT) and reinforcement learning (RL) optimization, T2I models exhibit substantial variance across different samples [tian2025unigen, verdun2025soft, li2025reflect]. To reduce this randomness, the agent generates candidate images and employs the unified multimodal model to evaluate their visual and semantic alignment. The image with the highest alignment score is selected as the final output.
Termination (STOP).
@@ 3.2 Action Space
| Unified | GPT-4o | 0.81 | 0.71 | 0.89 | 0.83 | 0.79 | 0.74 | 0.80 |
| Vanilla | 0.44 | 0.55 | 0.68 | 0.44 | 0.60 | 0.39 | 0.52 | |
| Bagel | ImAgent | 0.63 | 0.63 | 0.72 | 0.59 | 0.69 | 0.53 | 0.63 |
| Vanilla | 0.30 | 0.37 | 0.49 | 0.36 | 0.42 | 0.26 | 0.35 | |
| Janus-Pro-7B | ImAgent | 0.44 | 0.46 | 0.55 | 0.45 | 0.50 | 0.27 | 0.44 |
| Idiom | Textual | Entity | Scientific | Overall | |||||||
| Types | Model | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. |
@@
| Unified | Gemini-2.0 | 52.4 | 87.8 | 73.0 | 83.3 | 67.0 | 94.3 | 66.7 | 89.3 | 64.8 | 88.7 |
| Vanilla | 30.2 | 85.7 | 36.6 | 68.4 | 45.0 | 94.7 | 54.4 | 87.5 | 41.6 | 84.1 | |
| Bagel | ImAgent | 37.7 | 90.0 | 54.2 | 79.1 | 52.6 | 96.6 | 61.2 | 90.3 | 51.4 | 89.0 |
| Vanilla | 25.5 | 78.0 | 37.2 | 70.9 | 38.5 | 87.6 | 44.9 | 77.8 | 36.5 | 78.6 | |
| Janus-Pro-7B | ImAgent | 27.9 | 86.0 | 35.3 | 68.8 | 40.7 | 89.7 | 51.2 | 84.2 | 38.8 | 82.2 |
4 Experiments
4.1 Experimental Setting
Models.
We build our ImAgent on two of the most recent and powerful unified models, Bagel [deng2025emerging] and Janus-Pro-7B [chen2025janus]. Both models possess versatile multimodal capabilities, including image understanding, image generation, which collectively form the foundation of our framework. For image generation, we leverage both Bagel and Janus-Pro-7B. However, since Janus-Pro-7B does not support image editing, we employ Bagel exclusively for the image editing tasks.
Benchmarks.
We evaluate ImAgent on multiple benchmarks, including 3 benchmarks on image generation and 4 benchmarks for image editing tasks. For image generation, we use R2I-Bench [chen2025r2i], which is designed to rigorously assess reasoning-driven T2I generation; WISE [niu2025wise], which covers six categories of image generation scenarios; and T2I-ReasonBench [sun2025t2i], which assesses the models’ reasoning ability in generative tasks. For image editing, we adopt GEdit-Bench [liu2504step1x], which contains both Chinese and English instruction-based editing tasks; RISEBench [zhao2025envisioning], which focuses on reasoning-informed visual editing across diverse reasoning types; KRIS-Bench [wu2025kris], which evaluates reasoning capabilities over factual, conceptual, and procedural knowledge; and ImgEdit-Bench [ye2025imgedit], which is used to evaluate image editing performance in terms of instruction adherence, editing quality, and detail preservation.
Baselines.
For image generation, we employ SDv1.5 [rombach2022high], SD3-Medium [rombach2022high], SDXL [podell2023sdxl], SD3.5-Medium [esser2024scaling], SD3.5-Large [esser2024scaling], PixArt-Alpha [chen2024pixart], FLUX.1-Dev [blackforest2024flux], Sana-1.5 [xie2025sana], Lumina-T2I [qin2025lumina], LLM4GENSD1.5 [liu2025llm4gen], ELLASD1.5 [hu2024ella], LlamaGen [sun2024autoregressive], DALL-E-3 [ma2025learning], gpt-image-1 [hurst2024gpt], Omnigen [xiao2025omnigen], FLUX.1-schnell [flux2024], andPlayground-v2.5 [li2024playground], as generation-based models. We also include Janus [wu2025janus], VILA-U [wu2024vila], Show-o [xie2024show], Janus-Pro-7B [chen2025janus], Emu3 [wang2024emu3], Lumina-Image 2.0 [qin2025lumina], show-o [xie2024show], GoT [fang2025got] and MetaQuery-XL [pan2025transfer] as unified multimodal models.
| Types | Model | Temporal | Causal | Spatial | Logical | Overall |
| Gemini-2.0-Flash-pre | 10.6 | 13.3 | 11.0 | 2.3 | 9.4 | |
| Gemini-2.0 | 8.2 | 15.5 | 23.0 | 4.7 | 13.3 | |
| Private | GPT-4o | 34.1 | 32.2 | 37.0 | 10.6 | 28.9 |
| EMU2 | 1.2 | 1.1 | 0.0 | 0.0 | 0.5 | |
| OmniGen | 1.2 | 1.0 | 0.0 | 1.2 | 0.8 | |
| Step1X-Edit | 0.0 | 2.2 | 2.0 | 3.5 | 1.9 | |
| Qwen-Image-Edit | 4.7 | 10.0 | 17.0 | 2.4 | 8.9 | |
| FLUX.1-Kontext-Dev | 2.3 | 5.5 | 13.0 | 1.2 | 5.8 | |
| Ovis-U1 | 1.2 | 3.3 | 4.0 | 2.4 | 2.8 | |
| Open | Seedream-4.0 | 12.9 | 12.2 | 11.0 | 7.1 | 10.8 |
| Vanilla | 2.4 | 5.6 | 14.0 | 1.2 | 6.1 | |
| Bagel | ImAgent | 17.6 | 15.6 | 16.0 | 2.4 | 13.1 |
For image editing, we evaluate both private and open baselines. The private baselines include Gemini-2.0 [google2025_gemini2_image], Doubao [shi2411seededit], GPT-4o [openai2025gpt4o_image], and Gemini-2.0-Flash-pre [team2023gemini]. The open baselines include Step1X-Edit [liu2025step1x], Instruct-Pix2Pix [brooks2023instructpix2pix], MagicBrush [zhang2023magicbrush], AnyEdit [yu2025anyedit], OmniGen [xiao2025omnigen], EMU2 [sun2024generative], Qwen-Image-Edit [wu2025qwenimagetechnicalreport], FLUX.1-Kontext-Dev [labs2025flux1kontextflowmatching], Ovis-U1 [wang2025ovisu1], and Seedream-4.0 [seedream2509seedream].
4.2 Quantitative Results
Image Generation.
As shown in Table 2, ImAgent achieves outstanding performance on the WISE benchmark. Compared to their respective backbone models, ImAgent built upon Bagel and Janus-Pro-7B consistently surpasses the vanilla counterparts, with performance improvements of 21.2% and 25.7%, respectively, demonstrating the effectiveness of our approach in enhancing image generation quality through coordinated reasoning and refinement. Notably, the vanilla Janus-Pro-7B underperforms compared to SDXL and EMU3 and performs on par with Show-o. In contrast, ImAgent constructed on Janus-Pro-7B surpasses all of them, highlighting its strong test-time scaling capability and generalization potential. Furthermore, on the reasoning benchmarks, ImAgent consistently delivers strong results. As shown in Tables 1 and 3, ImAgent outperforms the vanilla models across both backbone architectures. For instance, ImAgent achieves improvements of 14.8% and 7.5% across both backbones on R2I-Bench, respectively. Moreover, ImAgent built upon Bagel surpasses the commercial model Gemini-2.0 in overall quality on the T2I-ReasonBench, further demonstrating its effectiveness in reasoning-based image generation.
Image Editing.
Furthermore, as shown in Figure 4, ImAgent also achieves excellent performance on KRISBench, attaining an overall score of 67.13, surpassing vanilla Bagel which scores 63.16. ImAgent performs well on ImgEdit-Bench (shown in Figure 3) as well, with vanilla achieving 2.89 and underperforming compared to Step1X-Edit, while ImAgent reaches 3.15, outperforming Step1X-Edit and achieving the best performance among open-source models. This further demonstrates the effectiveness of ImAgent in test-time scaling.
Moreover, ImAgent outperforms vanilla Bagel across both English and Chinese settings in GEdit-Bench, achieving improvements of 5.5% and 5.2%, respectively. Detailed results are provided in the Appendix.
4.3 Qualitative Analysis
Image Refinement Flow.
As shown in Figure 1, we present several qualitative examples, including two for image generation and two for image editing, each executed for five steps. From an overall perspective, we can observe that the image quality improves progressively with each step, demonstrating that ImAgent refines the image gradually.
Specifically, consider a case where the user aims to generate an image based on the prompt “moldy bread.” In the first step, ImAgent performs the Naive Generation action and produces an image containing fresh bread, which clearly deviates from the prompt. The policy controller then selects the Prompt Refinement action to improve the prompt using both the generated image and the user’s original input. Although the updated image shows slight improvement, it remains unsatisfactory. To address this, the controller chooses the Image Detail Refinement action, leading to the appearance of mold on the bread. Subsequently, ImAgent applies the Prompt Enhancement action to further strengthen the textual guidance, resulting in an image that better aligns with the user’s intent. Finally, recognizing that the prompt is now sufficiently clear but the visual result could still be improved, ImAgent executes the Best-of-N Sampling action to generate multiple candidates and select the most aligned one. The final output successfully captures the desired “moldy bread” concept.
Detailed Policy Controller Decision.
We present a detailed case study of the policy controller’s output to illustrate how it makes action decisions. As shown in Figure 6, which visualizes the decision-making process for the case discussed earlier, the policy controller selects the promptrefinement action to enrich the textual prompt and improve semantic clarity for subsequent generation steps. We present more examples in the Appendix.
4.4 Ablation Studies
Comparison with Self-CoT.
Since the vanilla Bagel model already incorporates self-CoT capabilities, we further compare ImAgent with self-CoT to evaluate the effectiveness of our approach. For the image editing task, as shown in Figure 5(a), ImAgent significantly outperforms both the vanilla and self-CoT variants on the Gedit-Bench across both English and Chinese settings. Furthermore, ImAgent also surpasses self-CoT in the image generation task. As illustrated in Figures 5(b) and (c), under the Janus-Pro-7B backbone, the performance of self-CoT notably drops compared to the vanilla model, whereas ImAgent consistently achieves superior results. Similarly, under the Bagel backbone, ImAgent maintains clear advantages over both the vanilla and self-CoT models. These findings suggest that the improvements brought by ImAgent stem not only from self-CoT–like reasoning refinement, but also from the different actions coordination enabled by the strong policy controller.
Effectiveness of the Whole ImAgent.
To evaluate the effectiveness of ImAgent in coordinately selecting and executing appropriate actions, we compare its performance with that of single-turn actions and a random action selection policy, where the for random selection is set the same as in ImAgent. As shown in Table 5, experiments are conducted on R2I-Bench [chen2025r2i]. All single-turn actions perform well compared to the naive generation action, which serves as the lower bound (0.54). Moreover, the random action selection policy achieves a score of 0.59, higher than all single-turn actions, indicating that naively constructing an agent is beneficial. Finally, ImAgent outperforms the random policy with a score of 0.62, demonstrating that it effectively and efficiently boosts test-time scaling in image generation.
| Method | Random | ImAgent | |||||
| Score | 0.54 | 0.58 | 0.57 | 0.56 | 0.56 | 0.59 | 0.62 |
5 Conclusion
Supplementary Material
Appendix A Algorithm for Image Editing
Appendix B Detailed Action Description
Naive Generation/Editing.
Prompt Enhancement with CoT.
| Method | Factual | Conceptual | Procedural | Overall | |||
| AP | SP | SS | NS | LP | ID | ||
| Vanilla | 73.27 | 67.17 | 61.10 | 62.36 | 54.38 | 64.77 | 63.16 |
| ImAgent | 74.39 | 71.15 | 69.80 | 66.03 | 54.83 | 68.78 | 67.13 |

### v3

@@ 1 Introduction
Text-to-Image (T2I) models [36, 35, 33, 44, 47, 34] have made remarkable progress in generating visually realistic and semantically coherent images from natural language descriptions. Despite these advancements, the quality of generated results often exhibits randomness and inconsistency with the given prompts [62, 18]. This limitation primarily stems from the strong dependency of existing T2I models on the clarity and specificity of textual inputs. When the prompt is vague or underspecified, the models tend to overlook key semantic elements, leading to images that deviate from the intended meaning and fail to fully capture the user’s intent.
To mitigate this problem, prior research has proposed various strategies, such as prompt rewriting [1, 29, 25], best-of-N sampling [21, 42], classifier-free guidance [11, 37, 39], and self-revision or iterative refinement [41, 19, 14]. These approaches aim to either clarify the textual prompt or reduce the model’s inherent randomness rather than finetuning the models, reflecting the underlying principle of test-time scaling—improving generation quality through additional inference-time computation.
@@
In this paper, we aim to construct a universal agent for image generation that can adaptively select the optimal action for a given case, allocate computational resources accordingly, and execute the selected action within the agent itself without relying on any external models. This design enables more efficient test-time scaling. Thanks to recent advances in unified multimodal models [6, 55, 55, 59, 34], which integrate text generation, image generation, and visual understanding within a single framework, this assumption becomes attainable. Building upon this foundation, we introduce ImAgent, a training-free unified multimodal agent designed to perform efficient test-time scaling for image generation.
@@ 2.1 Text-to-Image Generation
Text-to-Image (T2I) models have demonstrated remarkable ability in synthesizing visually realistic and semantically coherent images from natural language descriptions. The evolution of T2I models has followed several major paradigms, including diffusion-based models [33, 36, 35, 7, 9, 58], autoregressive models [44, 69, 47, 52, 73, 61, 60], and more recently, unified multimodal models [57, 59, 74, 49, 6, 72, 26]. These approaches have significantly advanced image fidelity, diversity, and text-image alignment.
@@ 2.2 Unified Multimodal Models
Unified multimodal models refer to models that jointly possess multimodal understanding and generation capabilities within a single framework. Unlike traditional pipelines that rely on separate models for text or image understanding, image generation, and image editing, these models integrate vision and language processing into a shared architecture, enabling seamless information exchange across modalities [57, 74, 49, 55, 55]. This convergence of understanding and generation not only improves efficiency and coherence but also paves the way for building more general-purpose, human-like agents capable of reasoning, creating, and self-correcting.
@@ 2.3 Generation Optimization Strategies
Despite the remarkable progress of T2I models, the quality of generated images still depends heavily on the clarity of prompts and the randomness inherent in the generation process. To alleviate these issues, various optimization strategies have been proposed. Best-of-N sampling [48, 50, 21] generates multiple candidates and selects the best one, reducing the influence of stochastic variations in model sampling. Prompt rewriting [48, 53, 64] reformulates or enriches the input text to provide more explicit guidance for the generator, improving semantic alignment and visual coherence. classifier-free guidance [11, 37, 39] controls the conditioning strength to balance fidelity and diversity, and iterative self-refinement [63, 28, 41, 19, 14] enables models to assess and revise their outputs through multiple rounds of generation. Although these techniques significantly enhance performance, they typically operate as independent modules with limited interaction between reasoning, generation, and evaluation—highlighting the need for unified multimodal frameworks that integrate these capabilities into a cohesive pipeline.
@@ 3 Method
We construct ImAgent upon unified multimodal models, which inherently possess both understanding and generation capabilities. Motivated by the observation that multimodal understanding is generally more reliable and easier to achieve than multimodal generation [70, 67], ImAgent leverages the model’s strong understanding ability to adaptively select the optimal action for a given case that enhances the overall generation quality. This design enables ImAgent to fully exploit the potential of a single unified multimodal model without any additional training or external modules, thereby achieving efficient test-time scaling. In this section, we first introduce the overall agent framework of ImAgent, followed by a detailed description of its action space.
@@ 3.1 Agent Framework Overview
This iterative process continues until the agent selects the STOP action, which indicates that the generated result is satisfactory, or until the maximum number of steps is reached. Notably, both the policy controller and the action execution are implemented within a single unified multimodal model. The detailed algorithm for image generation is presented in Algorithm 1, while the algorithm for image editing follows a similar procedure and is provided in the Appendix.
3.2 Action Space
In this section, we introduce the action space of ImAgent. The detailed definitions and usage of each action are provided in the Appendix 7. The detailed prompts for each action are provided in the Appendix 9.
Naive Generation/Editing.
Prompt Enhancement with CoT.
This action refines the input prompt by enriching vague or underspecified user queries with additional contextual and descriptive details. Since text-to-image (T2I) models are highly sensitive to prompt wording [1, 29, 25], more elaborate and specific prompts typically lead to higher-quality visual outputs. However, most T2I backbones are trained on simple captions or surface-level image descriptions, which limits their ability to reason over complex or compositional instructions. To address this limitation, we leverage the language reasoning capability of the model’s understanding module through Chain-of-Thought (CoT) prompting. By explicitly performing intermediate reasoning and elaboration, the agent transfers structured linguistic insights from the understanding domain to the generation process, thus producing more semantically aligned and visually coherent outputs.
Prompt Revision Based on the Generated/Edited Image.
Image Detail Refinement.
This action targets minor visual imperfections when the input instruction is already satisfactory and the remaining issues mainly originate from the generation module. It focuses on refining local visual details such as textures, lighting, and small artifacts. Specifically, the policy controller automatically generates a local editing prompt conditioned on the current image and the original user prompt, which is used internally to guide fine-grained image editing. Note that this local editing prompt is different from the user-provided prompt in image editing tasks. By enhancing fine-grained visual fidelity while preserving semantic consistency, this action improves the perceptual quality and realism of the generated results.
Best-of-N Sampling.
This action mitigates the inherent stochasticity of T2I generation. Unlike language models that typically produce stable, low-entropy outputs through supervised fine-tuning (SFT) and reinforcement learning (RL) optimization, T2I models exhibit substantial variance across different samples [48, 50, 21]. To reduce this randomness, this action generates candidate images and then evaluate their visual and semantic alignment based on the given prompt. The image with the highest alignment score is selected as the final output.
STOP.
3.3 Policy Controller
Decision Prompt.
We construct a structured decision prompt for the policy controller, which takes the current prompt, the generated image, and the observation history as inputs, and outputs a discrete action from the predefined action space . As shown in Figure 3, the decision prompt includes the current generation prompt and the observation history (including action and prompt histories), and explicitly instructs the model to select the most suitable next action from the predefined action space.
Action Parsing.
[…]
To mitigate hallucinated or invalid action outputs, we adopt a fallback strategy: if the parsed action is not in the predefined action set, we default to the Naive Generation/Editing action. This design ensures stable execution even when the policy controller produces malformed or out-of-distribution outputs. We provide a comprehensive robustness analysis of ImAgent in Section 4.3.
4 Experiments
4.1 Experimental Setting
@@
| Bagel | Vanilla | 0.47 | 0.65 | 0.53 | 0.68 | 0.40 | 0.54 |
| Self-Rewrite | 0.54 | 0.67 | 0.54 | 0.69 | 0.47 | 0.58 | |
| ImAgent | 0.58 | 0.68 | 0.62 | 0.71 | 0.53 | 0.62 | |
| Janus-Pro-7B | Vanilla | 0.45 | 0.60 | 0.46 | 0.46 | 0.36 | 0.47 |
| Self-Rewrite | 0.40 | 0.48 | 0.35 | 0.39 | 0.28 | 0.38 | |
| ImAgent | 0.47 | 0.58 | 0.52 | 0.54 | 0.44 | 0.51 |
| Types | Model | Temporal | Causal | Spatial | Logical | Overall |
| Private | Gemini-2.0-Flash-pre | 10.6 | 13.3 | 11.0 | 2.3 | 9.4 |
| Gemini-2.0 | 8.2 | 15.5 | 23.0 | 4.7 | 13.3 | |
| GPT-4o | 34.1 | 32.2 | 37.0 | 10.6 | 28.9 | |
| Open | EMU2 | 1.2 | 1.1 | 0.0 | 0.0 | 0.5 |
| OmniGen | 1.2 | 1.0 | 0.0 | 1.2 | 0.8 | |
| Step1X-Edit | 0.0 | 2.2 | 2.0 | 3.5 | 1.9 | |
| Qwen-Image-Edit | 4.7 | 10.0 | 17.0 | 2.4 | 8.9 | |
| FLUX.1-Kontext-Dev | 2.3 | 5.5 | 13.0 | 1.2 | 5.8 | |
| Ovis-U1 | 1.2 | 3.3 | 4.0 | 2.4 | 2.8 | |
| Seedream-4.0 | 12.9 | 12.2 | 11.0 | 7.1 | 10.8 | |
| Bagel | Vanilla | 2.4 | 5.6 | 14.0 | 1.2 | 6.1 |
| Self-Rewrite | 5.9 | 17.8 | 21.0 | 1.2 | 11.9 | |
| ImAgent | 17.6 | 15.6 | 16.0 | 2.4 | 13.1 |
Agentic Details.
The maximum number of steps is set to 5 by default, meaning that the agent will automatically select the STOP action if this limit is reached. For the Best-of-N Sampling action, we set by default.
Backbone Models.
We build our ImAgent on two of the most recent and powerful unified models, Bagel [6] and Janus-Pro-7B [5]. Both models possess versatile multimodal capabilities, including image understanding, image generation, which collectively form the foundation of our framework. For image generation, we leverage both Bagel and Janus-Pro-7B. However, since Janus-Pro-7B does not support image editing, we employ Bagel exclusively for the image editing tasks.
Benchmarks.
We evaluate ImAgent on multiple benchmarks, including 3 benchmarks on image generation and 3 benchmarks for image editing tasks. For image generation, we use R2I-Bench [4], which is designed to rigorously assess reasoning-driven T2I generation; WISE [30], which covers six categories of image generation scenarios; and T2I-ReasonBench [43], which assesses the models’ reasoning ability in generative tasks. For image editing, we adopt GEdit-Bench [23], which contains both Chinese and English instruction-based editing tasks; RISEBench [71], which focuses on reasoning-informed visual editing across diverse reasoning types; and ImgEdit-Bench [65], which is used to evaluate image editing performance in terms of instruction adherence, editing quality, and detail preservation.
Baselines.
For image generation, we employ SDv1.5 [36], SD3-Medium [36], SDXL [33], SD3.5-Medium [7], SD3.5-Large [7], PixArt-Alpha [3], FLUX.1-Dev [16], Sana-1.5 [58], Lumina-T2I [34], LLM4GENSD1.5 [22], ELLASD1.5 [12], LlamaGen [44], DALL-E-3 [27], gpt-image-1 [13], Omnigen [57], FLUX.1-schnell [17], andPlayground-v2.5 [20], as generation-based models. We also include Janus [55], VILA-U [56], Show-o [59], Janus-Pro-7B [5], Emu3 [52], Lumina-Image 2.0 [34], show-o [59], GoT [8] and MetaQuery-XL [32] as unified multimodal models.
For image editing, we evaluate both private and open baselines. The private baselines include Gemini-2.0 [10], Doubao [40], GPT-4o [31], and Gemini-2.0-Flash-pre [46]. The open baselines include Step1X-Edit [24], Instruct-Pix2Pix [2], MagicBrush [68], AnyEdit [66], OmniGen [57], EMU2 [45], Qwen-Image-Edit [54], FLUX.1-Kontext-Dev [15], Ovis-U1 [51], and Seedream-4.0 [38].
Meanwhile, we also compare ImAgent with the backbone unified model and a prompt self-rewrite baseline implemented using the same backbone model.
@@
| GPT-4o | 0.81 | 0.71 | 0.89 | 0.83 | 0.79 | 0.74 | 0.80 | |
| Bagel | Vanilla | 0.44 | 0.55 | 0.68 | 0.44 | 0.60 | 0.39 | 0.52 |
| Self-Rewrite | 0.76 | 0.69 | 0.75 | 0.65 | 0.75 | 0.58 | 0.70 | |
| ImAgent | 0.63 | 0.63 | 0.72 | 0.59 | 0.69 | 0.53 | 0.63 | |
| Janus-Pro-7B | Vanilla | 0.30 | 0.37 | 0.49 | 0.36 | 0.42 | 0.26 | 0.35 |
| Self-Rewrite | 0.21 | 0.24 | 0.26 | 0.18 | 0.20 | 0.16 | 0.21 | |
| ImAgent | 0.44 | 0.46 | 0.55 | 0.45 | 0.50 | 0.27 | 0.44 |
| Types | Model | Idiom | Textual | Entity | Scientific | Overall | |||||
| Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | ||
@@
| Gemini-2.0 | 52.4 | 87.8 | 73.0 | 83.3 | 67.0 | 94.3 | 66.7 | 89.3 | 64.8 | 88.7 | |
| Bagel | Vanilla | 30.2 | 85.7 | 36.6 | 68.4 | 45.0 | 94.7 | 54.4 | 87.5 | 41.6 | 84.1 |
| Self-Rewrite | 44.6 | 84.3 | 44.0 | 73.7 | 52.4 | 91.6 | 57.7 | 88.3 | 49.7 | 84.5 | |
| ImAgent | 37.7 | 90.0 | 54.2 | 79.1 | 52.6 | 96.6 | 61.2 | 90.3 | 51.4 | 89.0 | |
| Janus-Pro-7B | Vanilla | 25.5 | 78.0 | 37.2 | 70.9 | 38.5 | 87.6 | 44.9 | 77.8 | 36.5 | 78.6 |
| Self-Rewrite | 11.9 | 74.6 | 9.2 | 49.3 | 12.0 | 74.3 | 14.0 | 59.3 | 11.1 | 63.3 | |
| ImAgent | 27.9 | 86.0 | 35.3 | 68.8 | 40.7 | 89.7 | 51.2 | 84.2 | 38.8 | 82.2 | |